which of the following statements is (are) correct?
(x) Average variable cost equals variable costs divided by quantity produced.
(y) Marginal cost equals the change in variable costs divided by the change in quantity produced.
(z) If you find the difference between average total cost and average variable cost and then multiply that value by the specified quantity then the resulting value is the amount of fixed costs.
A. (x), (y) and (z) B. (x) and (y) only
C. (x) and (z) only D. (y) and (z) only
E. (x) only

At Annie’s Apple Pies, the total cost of producing 800 apple pies is $6,400. The variable cost of producing 800 apple pies is equal to $4,800. At Annie’s Apple Pies, the average fixed cost of the 800th apple pie is

A. more than $3.00
B. $3.00
C. $2.50.
D. $2.00.
E. less than $2.00

At Big Bert’s Basketballers, the total cost of producing 100 premium grade basketballs is $4,800. The marginal cost of producing an additional ten basketballs is $590. The average total cost of 110 basketballs
A. is $48. B. is $49.
C. is $59. D. is more than $59
E. None of the above
Answer #1

Ans 1.) Option C

Average variable Cost = (Total variable cost)/(Quantity)

AFC = average fixed cost = ATC - AVC

AFC = (Total fixed cost)/quantity

Total fuxed cost = (Average fixed cost)*(quantity)

The marginal cost refers to the change in the total cost when an extra unit of output is produced.

Mathematically,

Marginal cost = (change in total cost)/(change in output)

  

Answer 2) OPTION - D

Total Cost = $6400

Variable cost = $4800

Fixed Cost = 6400 - 4800 = $1600

Average fixed cost = 1600/800 = $2

Thus, average fixed cost of 800th apple pie is equal to $2.

Answer 3.) Option B

As marginal cost of additional 10 basketballs is equal to $590, so, we add this cost to the total cost of producing 100 basketballs.

Total cost of producing 110 basket balls = 4800 + 590 = $5390

Average cost = 5390/110 = $49